OpenStudy (anonymous):

The difference of twice a number and five is three. Find the number. Translate the word problem to an equation. Which steps describe how to solve the equation? add 2 then multiply by 1/5 multiply by 1/2 then add 5 add 5 then multiply by 1/2 subtract 5 then multiply by - 1/2

OpenStudy (anonymous):

2x-5=3

OpenStudy (anonymous):

Move -5 to the RHS. 2x = 8 Divide each side by 2. x = 4

OpenStudy (texaschic101):

twice a number = 2x... the difference means a subtraction problem... is means equals... the difference of twice a number and 5 is 3... 2x - 5 = 3 --- add 5 to both sides 2x = 8 -- divide both sides by 2, or you can multiply both sides by 1/2 x = 4 questions ?